Find all numbers whose sum when added to its multiplicative inverse is 13/7
Answer by stanbon(75887)

Find all numbers whose sum when added to its multiplicative inverse is 13/7
----
Solve:
x + 1/x = 13/7
7x^2 + 7 = 13x
----------
7x^2 -13x + 7 = 0
----------
x = [13 +- sqrt(169-4*7*7)]/14
----
x = [13 +- sqrt(-27)]/14
---
x is not a Real Number.
Cheers,
Stan H.